# Cross-Modality Perturbation Synergy Attack for Person Re-identification

### v2

@@ 4.0.3 Ablation Study
Our method is implemented based on UAP-Retrieval Li et al. (2019). To validate the effectiveness of the proposed method, we conducted experiments by adding augmentation (Cross-Modality Attack Augmentation) and CMPS to the baseline. Results with AGW baseline model are reported in Tab. 8. The No.1 line represents the UAP-Retrieval algorithm. In the table, ’Aug’ indicates the use of the cross-modality attack augmentation proposed in this paper.
@@
Table 5: Comparison with state-of-the-art methods on the SYSU dataset (Visible to Infrared).
Evaluation Before LTA Col.+Del. CMPS
Metrics Attack (CVPR2022) (TPAMI2023) (Ours)
$r=1$ 47.50 8.42 3.23 1.11
$r=10$ 84.39 21.25 14.48 8.67
$r=20$ 92.14 27.98 20.15 16.14
mAP 47.65 9.16 3.27 1.41
Table 6: Comparison with state-of-the-art method on transferability experiments in SYSU (Visible to Infrared)
[…]
Col.+Del. Ours Col.+Del. Ours
$r=1$ 6.56 2.07 5.35 1.94
$r=10$ 18.42 10.39 16.11 9.46
$r=20$ 25.81 19.35 23.68 17.53
mAP 7.25 3.41 6.12 2.80
The effectiveness of augmentation method. Cross-modality attack augmentation method is a component of our method. Through comparisons between No.1, No.2, and No.4, we can observe the improvements brought by this augmentation method. For example, on the RegDB dataset, using this augmentation (No.5) helps the attack algorithm reduce mAP from 6.87% to 5.11%, which is 1.76% lower than when it is not used (No.1). Similarly, with the use of CMPS, the mAP is reduced from 3.98% to 3.46% (No.4), which is 0.52% lower than without augmentation (No.3). These results indicate that employing appropriate data augmentation for cross-modality ReID adversarial attacks can enhance the universality of learned cross-modality perturbations. If not specified, our CMPS method in the experiments defaults to using this augmentation technique.
Fig. 5 displays the experimental results with augmentation performed at different probabilities. It can be observed that when the probability value is around 20%, it achieves optimal effectiveness in assisting the attack. If not specified, a probability value of 20% for augmentation is used by default in experiments.
4.0.4 Comparison with State-of-the-Art
By comparing Tab. 5 and Tab. 7, it can be observed that LTA shows a significant decrease in attack effectiveness in the ’Infrared to Visible’ scenario. For instance, LTA goes from mAP=9.16% to mAP=15.24%. On the other hand, Col.+Del., as a universal perturbation attack, also benefits from learning on a large-scale samples, but its performance is still noticeably limited. From Table 6, it can be observed that our method achieves better performance than Col.+Del. in transfer attacks across different baselines in cross-modality ReID.
The conclusions drawn from the above experiments are as follows: 1) Obtaining perturbations through single-modality information in cross-modality attacks often results in limited attack effectiveness. 2) The experiments with Col.+Del. indicate that universal perturbations are suitable for cross-modality attacks. However, the lack of consideration for modality differences can limit its attack performance. 3) Our method successfully narrows the gap between different modalities, more effectively capturing shared features across modalities.
Table 7: Comparison with state-of-the-art methods on the SYSU dataset (Infrared to Visible).
Evaluation Before LTA Col.+Del. CMPS
Metrics Attack (CVPR2022) (TPAMI2023) (Ours)
$r=1$ 54.17 20.96 4.12 1.45
$r=10$ 91.14 32.18 16.85 7.71
$r=20$ 95.98 36.80 21.27 10.72
mAP 62.97 15.24 3.89 1.25
@@ 5 Conclusion
In this study, we have introduced a novel cross-modality attack method known as Cross-Modality Perturbation Synergy (CMPS) attack, aimed at evaluating the security of cross-modality ReID systems.
The core idea behind CMPS attack is to capture shared features between visible and non-visible images to optimize perturbations. Additionally, to complement CMPS, we have introduced targeted data augmentation techniques by converting images into grayscale images, further enhancing the attack performance. Through experiments conducted on the RegDB and SYSU datasets, we demonstrated the effectiveness of the proposed method while also revealing the limitations of traditional attack approaches.
@@
Our primary objective has been to assess the security of cross-modality ReID systems. In future research endeavors, we plan to develop robust ReID methods tailored specifically for cross-modality attacks, aimed at defending against adversarial samples. This study not only contributes to advancing the understanding of the security of cross-modality ReID systems but also provides a strong motivation for ensuring the reliability and security of these systems in real-world applications.

### v3

@@ 4.3 Ablation Study
Our method is implemented based on UAP-Retrieval [16]. To validate the effectiveness of the proposed method, we conducted experiments by adding augmentation (Cross-Modality Attack Augmentation) and CMPS to the baseline. Results with AGW baseline model are reported in Tab. 5. The No.1 line represents the UAP-Retrieval algorithm. In the table, ’Aug’ indicates the use of the cross-modality attack augmentation proposed in this paper.
@@
The effectiveness of augmentation method. Our approach includes cross-modality attack augmentation. Comparing results of No.1, No.2, and No.4 shows its benefits. For example, on the RegDB dataset, augmentation (No.2) reduces mAP from 6.87% to 5.11%, 1.76% lower than without augmentation (No.1). Similarly, with CMPS, mAP drops from 3.98% to 3.46% (No.4), a 0.52% decrease compared to No.3. These findings suggest that using appropriate augmentation enhances cross-modality ReID adversarial attacks’ universality. If not specified, our experiments default to using CMPS augmentation. Fig. 5 displays the experimental results with our augmentation performed at different probabilities. It can be observed that when the probability value is around 20%, it achieves optimal effectiveness in assisting the attack. If not specified, a probability value of 20% for augmentation is used by default in experiments.
Impact of adversarial boundary size. We conducted an ablation study on different adversarial boundary sizes ($\epsilon$), as shown in the supplementary material  6. In practical applications, $\epsilon$ is typically kept moderate to balance perturbation visibility and attack effectiveness. To maintain consistency with previous work  [39], we set $\epsilon=8$ for comparison unless otherwise specified.
@@ 5 Conclusion
In this study, we have proposed a cross-modality attack method known as Cross-Modality Perturbation Synergy (CMPS) attack, aimed at evaluating the security of cross-modality ReID systems. The core idea behind the CMPS attack is to capture shared knowledge between visible and non-visible images to optimize perturbations. Additionally, we proposed a cross-modality attack augmentation method, utilizing grayscale images to bridge the gap between different modalities, further enhancing the attack performance. Through experiments conducted on the RegDB and SYSU datasets, we demonstrated the effectiveness of the proposed method while also revealing the limitations of traditional attack approaches. The primary objective of this study has been to assess the security of cross-modality ReID systems. In future research, we plan to develop robust ReID methods specifically tailored for cross-modality attacks, aimed at defending against adversarial samples. This study not only contributes to advancing the understanding of the security of cross-modality ReID systems but also provides strong motivation for ensuring the reliability and security of these systems in real-world applications.
Acknowledgment
This work was supported in part by the National Natural Science Foundation of China under Grant No. 62276222 and the Public Technology Service Platform Project of Xiamen City, Grant No. 3502Z20231043.
Supplemental Material
6 Supplemental Experiments
Figure 6: Transferability experiments of the proposed method across different models on the RegDB dataset (visible to thermal). Transferability experiments of the proposed method across different models on the SYSU dataset (visible to Infrared).
Figure 7: Comparison of Transferability Between Different Methods on Two Cross-Modal Datasets SYSU and RegDB.
Table 6: Using the AGW baseline on the RegDB dataset, we conduct an ablation study to evaluate the impact of the adversarial boundary $\epsilon$ on the effectiveness of the proposed CMPS attack (rank-1 accuracy).
Adversarial Boundary ($\epsilon$) Visible to Thermal Thermal to Visible
- 70.0% 70.5%
2 32.7% 40.5%
4 9.6% 13.8%
8 2.3% 2.0%
16 0.3% 0.5%
7 Proof of Method Superioritys
We design a cross-modality triplet loss to simultaneously optimize two modalities, which effectively captures common features between different modalities and enhances the cross-modality adaptability of universal perturbations.
7.1 Definition of Cross-Modality Triplet Loss
The cross-modality triplet loss aims to optimize the model by adjusting the distance relationships among triplet samples (anchor, positive, negative) so that samples of the same identity are closer, while samples of different identities are farther apart. Specifically, given samples $(x_{\text{A}},x_{\text{P}},x_{\text{N}})$, where:
  • $x_{\text{A}}$ is the anchor sample,
  • $x_{\text{P}}$ is the positive sample with the same identity as the anchor (from a different modality),
  • $x_{\text{N}}$ is the negative sample with a different identity from the anchor.
The triplet loss function is defined as:
$L_{\text{triplet}}=\max\left(0,D(f(x_{\text{A}}),f(x_{\text{P}}))-D(f(x_{\text{A}}),f(x_{\text{N}}))+\alpha\right)$ (10)
where $D(\cdot,\cdot)$ denotes the distance metric (e.g., Euclidean distance), and $\alpha$ is a margin hyperparameter.
Mathematically, given the cross-modality triplet loss:
$L_{\text{triplet}}=\max\left(\left(\lVert C^{n}_{g}-f^{adv}_{RGB}\rVert_{2}-\lVert C^{p}_{ir}-f^{adv}_{RGB}\rVert_{2}+\rho\right),0\right)$ (11)
We can view it as part of the sum of the loss functions for two modalities:
$\mathcal{L}_{A}(\eta)=\lVert C^{n}_{g}-f^{adv}_{RGB}\rVert_{2}$ (12)
$\mathcal{L}_{B}(\eta)=\lVert C^{p}_{ir}-f^{adv}_{RGB}\rVert_{2}$ (13)
Thus, the overall optimization objective can be expressed as:
$\eta_{\text{agg}}^{*}=\arg\min_{\eta}\left(\mathcal{L}_{A}(\eta)+\mathcal{L}_{B}(\eta)+\rho\right)$ (14)
This form effectively aggregates the losses of different modalities, thereby optimizing the loss functions of different modalities simultaneously, achieving joint optimization of cross-modality data. This approach trains universal perturbations with better generalization capabilities than methods that consider only single-modality information.
7.2 Proof of Aggregated Optimization Superiority
Assume we have data from two modalities: modality A and modality B. Let $\mathcal{L}_{A}(\eta)$ and $\mathcal{L}_{B}(\eta)$ be the loss functions on modality A and modality B, respectively. The objective of single-modality training is:
$\min_{\eta}\mathcal{L}_{A}(\eta)+\mathcal{L}_{B}(\eta)$ (15)
The stepwise optimization method first optimizes $\mathcal{L}_{A}(\eta)$ and then optimizes $\mathcal{L}_{B}(\eta)$:
$\eta^{*}=\arg\min_{\eta}\mathcal{L}_{A}(\eta)\rightarrow\eta^{**}=\arg\min_{\eta}\mathcal{L}_{B}(\eta^{*})$ (16)
The aggregated optimization of the two loss functions is:
$\eta_{\text{agg}}^{*}=\arg\min_{\eta}\left(\mathcal{L}_{A}(\eta)+\mathcal{L}_{B}(\eta)\right)$ (17)
Using the gradient aggregation method, it can be expressed as:
$\nabla_{\eta}\mathcal{L}_{\text{agg}}=\nabla_{\eta}\left(\mathcal{L}_{A}(\eta)+\mathcal{L}_{B}(\eta)\right)$ (18)
Next, we consider the different optimization paths of the two methods.
7.2.1 Stepwise Optimization Method
The stepwise optimization method first optimizes the loss function of modality A and then the loss function of modality B. Assume the update rule at iteration $k$ is:
$\eta^{(k+1)}=\eta^{(k)}-\alpha\nabla_{\eta}\mathcal{L}_{A}(\eta^{(k)})$ (19)
After optimizing the loss function of modality A, the loss function of modality B is optimized:
$\eta^{(k+1)}=\eta^{(k)}-\alpha\nabla_{\eta}\mathcal{L}_{B}(\eta^{(k)})$ (20)
Since the two optimization processes are separate, this may result in $\eta$ being optimal for modality A but not necessarily for modality B.
7.2.2 Aggregated Optimization Method
The aggregated optimization method considers the losses of both modalities in each iteration. Assume the update rule at iteration $k$ is:
$\eta^{(k+1)}=\eta^{(k)}-\alpha\left(\nabla_{\eta}\mathcal{L}_{A}(\eta^{(k)})+\nabla_{\eta}\mathcal{L}_{B}(\eta^{(k)})\right)$ (21)
In this way, each update considers the losses of both modalities, ensuring that $\eta$ approaches the optimal solution for both modalities.
To further prove that the aggregated optimization method can find a better perturbation $\eta$, we can analyze the existence and uniqueness of the optimal solution.
Assume $\mathcal{L}_{A}(\eta)$ and $\mathcal{L}_{B}(\eta)$ are continuously differentiable and convex loss functions. According to convex optimization theory, the optimal solutions of the loss functions exist and are unique.
The optimal solution of the stepwise optimization method is:
$\eta_{\text{step}}^{*}=\arg\min_{\eta}\left(\mathcal{L}_{A}(\eta)+\mathcal{L}_{B}(\eta^{*})\right)$ (22)
where $\eta^{*}$ is the optimal solution of $\mathcal{L}_{A}(\eta)$.
The optimal solution of the aggregated optimization method is:
$\eta_{\text{agg}}^{*}=\arg\min_{\eta}\left(\mathcal{L}_{A}(\eta)+\mathcal{L}_{B}(\eta)\right)$ (23)
Since $\eta_{\text{step}}^{*}$ is not necessarily globally optimal for modality B, and $\eta_{\text{agg}}^{*}$ is the global optimal solution considering both modalities, we can derive:
$\mathcal{L}_{A}(\eta_{\text{agg}}^{*})+\mathcal{L}_{B}(\eta_{\text{agg}}^{*})\leq\mathcal{L}_{A}(\eta_{\text{step}}^{*})+\mathcal{L}_{B}(\eta_{\text{step}}^{*})$ (24)
7.3 Generalization Error Analysis
Generalization error measures the model’s performance on unseen data. We can further prove the superiority of aggregated training through generalization error analysis.
Let $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ be the losses on the training and test sets, respectively. The generalization error is defined as:
$\mathcal{E}_{\text{gen}}=\mathcal{L}_{\text{test}}(\eta)-\mathcal{L}_{\text{train}}(\eta)$ (25)
The upper bound of the generalization error can be expressed using measures such as Rademacher complexity or VC dimension. For machine learning models, the lower the model complexity, the smaller the generalization error. Simultaneously optimizing the losses for multiple tasks (modalities) can reduce overfitting to a single task (modality), as the model needs to perform well on multiple tasks (modalities) simultaneously. This effectively introduces an implicit regularization effect, reducing the model complexity. Therefore, compared to the stepwise optimization method, the aggregated optimization method can effectively reduce the complexity of the perturbation model. The lower the model complexity, the smaller the generalization error.
The Rademacher complexity measures the complexity of a class of models on a given sample set. For a function $h$ in the hypothesis space $\mathcal{H}$, the empirical Rademacher complexity on $n$ samples is defined as:
^$\hat{\mathcal{R}}_{n}(\mathcal{H})=\mathbb{E}_{\sigma}\left[\sup_{h\in\mathcal{H}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}h(x_{i})\right]$ (26)
where $\sigma_{i}$ are Rademacher random variables, taking values ±$\pm 1$ with equal probability.
The impact of modality aggregation on complexity:
Assume $\mathcal{H}_{\text{A}}$ and $\mathcal{H}_{\text{B}}$ are the hypothesis spaces of modality A and modality B, respectively. The stepwise optimization method first optimizes $\mathcal{H}_{\text{A}}$ and then $\mathcal{H}_{\text{B}}$. Its empirical Rademacher complexity can be expressed as:
^$\hat{\mathcal{R}}_{n}(\mathcal{H}_{\text{step}})=\hat{\mathcal{R}}_{n}(\mathcal{H}_{\text{A}})+\hat{\mathcal{R}}_{n}(\mathcal{H}_{\text{B}})$ (27)
The aggregated optimization method optimizes $\mathcal{H}_{\text{A}}\cup\mathcal{H}_{\text{B}}$ simultaneously. Its empirical Rademacher complexity is:
^$\hat{\mathcal{R}}_{n}(\mathcal{H}_{\text{agg}})=\hat{\mathcal{R}}_{n}(\mathcal{H}_{\text{A}}\cup\mathcal{H}_{\text{B}})$ (28)
According to the properties of Rademacher complexity, the complexity of $\mathcal{H}_{\text{A}}\cup\mathcal{H}_{\text{B}}$ is usually less than or equal to the sum of the complexities of $\mathcal{H}_{\text{A}}$ and $\mathcal{H}_{\text{B}}$:
^$\hat{\mathcal{R}}_{n}(\mathcal{H}_{\text{agg}})\leq\hat{\mathcal{R}}_{n}(\mathcal{H}_{\text{A}})+\hat{\mathcal{R}}_{n}(\mathcal{H}_{\text{B}})$ (29)
Generalization error upper bound derivation:
Using Rademacher complexity, we can derive the upper bound of the generalization error. For the loss function $\mathcal{L}$ and hypothesis space $\mathcal{H}$, the upper bound of the generalization error is:
$\mathcal{E}_{\text{gen}}\leq 2\hat{\mathcal{R}}_{n}(\mathcal{L}\circ\mathcal{H})+\mathcal{O}\left(\frac{1}{\sqrt{n}}\right)$ (30)
where $\mathcal{L}\circ\mathcal{H}$ denotes the composition of the loss function with the hypothesis space.
The upper bound of the generalization error for the stepwise optimization method is:
$\mathcal{E}_{\text{gen, step}}\leq 2\left(\hat{\mathcal{R}}_{n}(\mathcal{L}\circ\mathcal{H}_{\text{A}})+\hat{\mathcal{R}}_{n}(\mathcal{L}\circ\mathcal{H}_{\text{B}})\right)+\mathcal{O}\left(\frac{1}{\sqrt{n}}\right)$ (31)
The upper bound of the generalization error for the aggregated optimization method is:
$\mathcal{E}_{\text{gen, agg}}\leq 2\hat{\mathcal{R}}_{n}(\mathcal{L}\circ(\mathcal{H}_{\text{A}}\cup\mathcal{H}_{\text{B}}))+\mathcal{O}\left(\frac{1}{\sqrt{n}}\right)$ (32)
Since
^$\hat{\mathcal{R}}_{n}(\mathcal{L}\circ(\mathcal{H}_{\text{A}}\cup\mathcal{H}_{\text{B}}))\leq\hat{\mathcal{R}}_{n}(\mathcal{L}\circ\mathcal{H}_{\text{A}})+\hat{\mathcal{R}}_{n}(\mathcal{L}\circ\mathcal{H}_{\text{B}})$ (33)
Therefore:
$\mathcal{E}_{\text{gen, agg}}\leq\mathcal{E}_{\text{gen, step}}$ (34)
This indicates that the aggregated optimization method has a lower upper bound on the generalization error compared to the stepwise optimization method.
8 Discussion
8.1 Ethical Considerations
In this study, we introduce a novel cross-modal adversarial attack method known as Cross-Modality Perturbation Synergy (CMPS). This research offers a new perspective on understanding and enhancing the security of cross-modal ReID systems by leveraging shared features across different modalities to optimize perturbations. However, this approach also raises a series of ethical and safety concerns regarding the potential negative impacts of adversarial attack techniques. The CMPS method, like other adversarial technologies, can be maliciously exploited, posing a serious threat to public safety.
However, we recognize the positive value of adversarial attack research. It reveals vulnerabilities in existing systems, prompting academia and industry to make in-depth improvements to the robustness of machine learning models. The positive impact of this study lies in its potential to combine adversarial training with the attack methods presented to enhance system security and bring positive social impacts. Therefore, we emphasize the importance of conducting adversarial attack research within an ethical framework and encourage further development of defensive technologies to build a safer and more reliable technological environment.
8.2 Limitations and Future Work
Here, we need to acknowledge the limitations of the proposed method and identify potential directions for future research. Firstly, current attack techniques primarily focus on gradient-based perturbation optimization for given datasets. However, in real-world scenarios, the modalities encountered are often unknown and not limited to RGB, infrared, and thermal imaging. Moreover, effectively transferring perturbations to different and unknown modalities presents a significant research challenge.
When dealing with various models and modalities, gradient-based methods face several challenges. Firstly, these methods are prone to "catastrophic forgetting," where learning new information can lead to the loss of previously learned knowledge, affecting the effectiveness of perturbations. Secondly, the inconsistency of gradient information across multiple models and modalities can negatively impact the stability and generalizability of the method. Therefore, future research should explore more robust algorithms that can effectively operate in complex environments involving multiple modalities and models, thereby enhancing the applicability and transferability of attacks.